OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3 0 −3 3 2

OpenStudy (anonymous):

use substitution y=x-3 4x+2(x-3)=12 4x+2x-6=12 +6 +6 _______________ 6x=18 x=3 substitute x in original equation 3-y=3 -3 -3 ______________ -y=0 y=0 (3,0)

OpenStudy (anonymous):